You planned to buy a stock of company A, given the following information of the stock:

i. Beta of the stock: 1.5

ii. 3 months treasury bill yield: 3%

iii. Market return: 12%


a.               Calculate your required rate of return for the stock?          


b.              Assuming the stock price paid RM 2 dividend last period, and expected to grow at 8 % for 4 years, and a constant rate of 4% after that, what is the expected stock price that you will buy the stock in order to earn your required return?          




1
Expert's answer
2022-01-11T11:19:32-0500

a.


"CARM=Rf+\beta(Rm-Rf)=3+1.5(12-3)=16.5"


b.

"P=\frac{D(1+g)}{1+r}+\frac{D(1+g)^2}{(1+r)^2}+\frac{D(1+g)^3}{(1+r)^3}+\frac{D(1+g)^4}{(1+r)^4}=8.79"